solve using the elimination method 3x+6y=3 2x+9y=-8
Answer by TimothyLamb(4379)

3x+6y=3
2x+9y=-8
---
put the system of linear equations into standard form
---
3x+6y=3
2x+9y=-8
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 5
y = -2
---
now by elimination:
---
-2(3x + 6y = 3)
3(2x + 9y = -8)
---
-6x - 12y = -6
6x + 27y = -24
---
add the two equations:
---
15y = -30
y = -2
---
3x + 6y = 3
3x + 6*(-2) = 3
3x = 3 + 12
3x = 15
x = 5
---
solution by elimination:
x = 5
y = -2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---